1) The grades of students in a class are distributed normally. The average grade of the students in this class is 65; standard deviation is 2. The students who take the teacher give more points by making "systematic error". What is the passing score since it gives more than 25% of the class score in total? Calculate.
Answer #1

Solutions Given data, celo o = 2 4 = 65 P(x >x) = 25% = 0.25 P(x-x) = 1-025 = 0,75 PCX LX) - P(Z < 0.6745 ) x = l + Zo x=65+